### v1

# Range-Limited Heaps’ Law for Functional DNA Words in the Human Genome

Abstract
Heaps’ or Herdan’s law is a linguistic law describing the relationship between the vocabulary/dictionary size (type) and word counts (token) to be a power-law function. Its existence in genomes with certain definition of DNA words is unclear partly because the dictionary size in genome could be much smaller than that in a human language. We define a DNA word in a genome as a DNA coding region that codes for a protein domain. Using human chromosomes and chromosome arms as individual samples, we establish the existence of Heaps’ law in the human genome within limited range. Our definition of words in a genomic or proteomic context is different from that in large language models for DNA or protein sequences where words are usually short. Although an approximate power-law distribution of protein domain sizes due to gene duplication and the related Zipf’s law is well known, their translation to the Heaps’ law in DNA words is not automatic. Several other animal genomes are shown herein also to exhibit range-limited Heaps’ law with our definition of DNA words, though with various exponents, partially depending on their level of complexity. Investigation of Heaps’ law and its exponent value could provide an alternative narrative of reusage and redundancy of protein domains as well as creation of new protein domains from a linguistic perspective.
Abbreviations: CRCPD: coding regions that code for protein domains.
1 Introduction
The study of molecular biology is full of examples in terminology, and in quantitative and empirical laws, borrowed from linguistics (Key, 2000). Jargons used in molecular biology, such as transcription, translation, code, etc. are often borrowed from terms in linguistics, not to mention the ubiquitous use of letters (nucleotides and amino acids) and text (DNA and protein sequences). The recent Nobel-prize-winning biotechnique of CRISPR was described as an equivalence of a text editor (Nelson et al., 2015). Not surprisingly, linguistic laws have also seen their mentioning in genomics. For example, Zipf’s law (Zipf, 1935) which describes how the word frequency in a text drops off with its rank, was tested in DNA sequences with k-mers/k-tuples (e.g. k=6) considered as “word” (Mantegna et al., 1994). This study was criticized, putting aside the issue of the definition of DNA words, because the rank-frequency plot did not show a good power-law as in the case of human languages (Konopka and Martindale, 1995; Li, 2002). Another example of linguistic law is the Menzerath’s law (Menzerath, 1928), which describes the tendency for using smaller linguistic units when these units are more numerous. The Menzerath’s law at the human exon level is the tendency for genes with more exons to have smaller exon sizes (Li, 2011a; Nikolaou, 2014), while other choices of units are also investigated (Semple et al., 2022).
In this study, we focus on another linguistic law, the Heaps law (Heaps, 1978) (or Herdan-Heaps law (Egghe, 2007)), which states that the number of unique words (vocabulary/dictionary size) as a function of text length is a power-law. Heaps’ law can be understood as a type(size)-token(size) (Herdan, 1960; Wetzel, 2009) relationship: type being a unique word, token being the usage of a word, i.e., the total number of types is the vocabulary size, and total number of tokens is the vocabulary usage. Heaps’ law can also be understood as a law of diminishing return. If one searches longer and longer texts for the purpose of finding new words, Heaps’ law indicates that the chance of finding new words becomes smaller and smaller as more texts are looked over. The latter meaning in Heaps law has been implied in the number of genes or gene-families discovered by sequencing more and more species’ genomes in a clade (Medini et al., 2020), or the number of single-nucleotide-polymorphisms (SNPs) discovered by sequencing more samples (Ionita-Laza et al., 2009).
In order to apply a linguistic law to genomics, one first needs to find an entity corresponding to a word. A common practice is to consider (overlapping) k-mers with a prefixed k value as “words” due to its easiness to calculate, its use in alignment-free sequence comparison (Zielezinski et al., 2017), and its importance in many bioinformatics algorithms (Chikhi and Medvedev, 2013; Koren et al., 2017; Rahman et al., 2018). However, the problem with k-mers as words is that k-mers lack functional meaning, although there are attempts to filter k-mers by other criteria and by relaxing the fixed k value to obtain some meaning (Bussemaker et al., 2000; Moghaddasi et al., 2017). In the framework of semiotics (Hoopes, 2011), word is part of the signifier-signified pair that represents a meaning. Without a biological meaning, a proposed DNA or protein word would be much less interesting if we continue to discuss linguistic laws. Although the in-frame non-overlapping 3-mers within exons (codons) do have a biological meaning, k-mers, generally speaking, do not.
Recently, large language models (Brant et al., 2007), such as bidirectional encoder representations from transformers (BERT) (Devlin et al., 2018) or generative pre-trained transformer (GPT) (Radford et al., 2019), have been applied to biomolecular sequences, in particular protein sequences (Rao et al., 2020; Madani et al., 2023). One aspect of natural language process is tokenization which means to delineate a character string by boundaries (Webster and Kit, 1992). For protein sequences, the tokenization is a process to find “words” (Ofer et al., 2021). The protein “words” found by this method are usually short, with average size of only a few (e.g. four) amino acids (Ferruz et al., 2022).
In this study, we use DNA segments that code a protein domain (Wang et al., 2021) as our DNA “words”. Treating protein domains as protein words has been proposed before (Searls, 2002; Gimona, 2006; Scaiewicz and Levitt, 2015; Yu et al., 2019; Buchan and Jones, 2020). Proposals for DNA words were usually limited to k-mers (Brendel et al., 1986) which according to our discussion above is not ideal. There were studies that consider gene internal structure in DNA sequences as “grammar” or rules (Done and Searls, 1994), but the emphasis of these were not on the definition of words. We prefer protein domain reverse-transcribed to DNA sequences as DNA words. By doing so, our definition of DNA words has a biological meaning, i.e. coding for a protein domain.
To investigate if Heaps’ law holds in the human genome, our task is to find all Coding Regions that Code for a Protein Domain (CRCPD). For this we need to count the number of times any CRCPD appears (number of tokens) and the number of different types of CRCPD appear (number of types). For one chromosome of the human genome, this pair of counts (#token, #type) will only give us one point. To see a relationship between token and type, we repeat the same counting process for each one of the 23 chromosomes and/or each one of the chromosome arms, so more points are produced. This may be chosen if more points are desired, as chromosomal arms may be seen as relatively independent entities, and different from the whole chromosomes where they belong.
It has been known for over twenty years that the histogram of protein domains or distinct parts follow an inverse power-law distribution in genomes (Qian et al., 2001; Luscombe et al., 2002; Koonin et al., 2002). Another well known property is the importance role that gene duplication plays in shaping the distribution of these protein domains as well as other genomic units (Müller et al., 2002; Gao and Miller, 2011; Li et al., 2016), echoing theoretical studies linking duplication and power-law behavior (Li, 1991; Ispolatov et al., 2005). A perfect inverse power-law histogram (distribution) will lead to a similar inverse power-law with rank as the x-variable, i.e., the Zipf’s law (Newman, 2005).
However, more and more studies show that when tokens are randomly chosen from a dictionary with a finite size whose words follow Zipf’s law distribution, the resulting type-token plot does not follow an exact Heaps’ law (Bernhardsson et al., 2009; Font-Clos and Corral, 2015). In particular, the type-token plot in log-log scale would show a concave downward shape (Font-Clos and Corral, 2015). Nevertheless, within a limited range, this concave curve might be approximately viewed as a straight line. Also, publications with proofs on relations between two power-law exponents in the two laws rely on certain assumptions, e.g. infinite dictionary size, and working in the infinite token limit (Boytsov, 2017). All these observations imply that the power-law distribution in protein domain known for the last twenty years does not automatically lead to a power-law type-token plot for CRCPD, which needs to be checked from the data directly.
In the next section, we describe the data and the methodology used. In section 3, we present evidence of the existence of range-limited Heaps’ law in the human genome; while we point out one chromosome (chromosome 19) as an outlier to the Heaps’ law; We also include in our study another level of genomic entities, chromosome arms, to explore the extent of Heaps’ law. In section 4, we examine the potential Heaps’ law in other animal genomes. In section 5, we generate artificial units by randomly sampling DNA word tokens from the list of known DNA word types, and confirm the previous results of log-log concave downward and further reconfirm a systematic deviation from the Heaps’ law. The same random sampling process is also carried out at the individual chromosome level, showing a distinction between the concavity and finite dictionary size effect. Finally, various points are discussed in the Discussion section, including a comparison between the creation of new words in human languages and evolution of protein domains in genomes.
2 Data and Methods
The protein domains used in the present study refer to Pfam (Mistry et al., 2021) (https://pfam.xfam.org/ or https://pfam-legacy.xfam.org/, which is hosted by InterPro (Paysam-Lafosse et al., 2022) after January 2023: https://www.ebi.ac.uk/interpro). The chromosomal locations of CRCPD in human genes are available from the University of California at Santa Cruz (UCSC) Genome Browser track: Pfam domains in GENECODE genes. A description of the methods used to derive this track can be found on UCSC Genome Browser site (https://genome.ucsc.edu/cgi-bin/hgTrackUi?db=hg19&g=ucscGenePfam).
To obtain this Pfam domains track in the human genome (for GRCh38/hg38), from the top bar of the genome browser, we select Tools $\rightarrow$ Table Browser, group = “Genes and Gene Predictions”, track = “Pfam in GENECODE”. The resulting table contains chromosome number, chromosomal location (start, end, in bp), name of the Pfam domain, etc.
In this study, we also used chromosome arms as units. To determine the chromosome arm to which a region belongs to, we download the cytoband information from:
https://hgdownload.cse.ucsc.edu/goldenpath/hg38/database/cytoBand.txt.gz to find the centromere locations in each chromosome. The counting of Pfam domains (both token and type) was carried out by R (https://www.r-project.org) scripts written by us.
The Heaps’ law is written as ($y$= vocabulary size = #type, or the number of unique CRCPD domains in a chromosome, $x$=vocabulary usage=#token, or number of times any of these CRCPD appear in the same chromosome):
$y=Cx^{\alpha}$ (1)
which after a logarithmic transformation on both side ($y^{\prime}=log(y)$, $x^{\prime}=log(x)$, $C^{\prime}=log(C)$): $y^{\prime}=C^{\prime}+\alpha x^{\prime}$ becomes a linear relationship. The exponent $\alpha$ is determined by a linear regression in R (the function lm). The AIC (for model selection) (Akaike, 1974; Li and Nyholt, 2001) calculation is carried out by the AICcmodavg package (cran.r-project.org/web/packages/AICcmodavg/), using the function aictab.
3 Results for the Human Genome
3.1 Heaps’ law for Pfam CRCPD holds true
The number of Pfam domain appearance in each autosomal and X chromosome in the human genome (version hg38) is obtained from the Pfam-in-GENECODE track of the UCSC genome browser. This would produce 23 points for checking the vocabulary size (type) and vocabulary usage (token) relationship and a regression. In order to increase the sample size, we also consider data from the p- and q-arm in a chromosome, as partitioned by the centromere for each metacentric (non-acrocentric) chromosome (41 points).
Fig.1(A) shows the number of types (unique number of Pfam domains) as a function of number of tokens (Pfam domain counts) for 23 human chromosomes in log-log scale. With the exception of one point, all other points scatter around a straightline, indicating a power-law and Heaps’ law. The outlier is chromosome 19, being lower in number of CRCPD types or higher in number of CRCPD tokens. The regression for log-token over log-type produces the slope (scaling exponent) $\alpha=0.658$ and $C=5.12$ in Eq.(1).
Fig.1(B) shows a similar type-token plot in log-log scale using chromosome arm data. The regression results are similar: $\alpha=0.683$ and $C=3.97$. Using both whole chromosome and chromosome arm data leads to Fig.1(C), with $\alpha=0.695$ and $C=3.7$. Figure 1(D) shows the same data as Fig.1(C) but in linear-linear scale. It is clear that power-law curve function fits the data better than a linear regression. In fact, the variance explained by regression is much higher in log-log scale ($R^{2}=0.887,0.873,0.9$ for chromosome data, chromosome arm data, and combined) than in linear-linear scale ($R^{2}=0.697,0.667,0.738$).
Removing the outliers (one point in Fig.1(A), two points in Fig.1(B) for two chromosome 19 arms, and three points in Fig. 1(C)) result in even better $R^{2}$: 0.968, 0.955, and 0.967. The slope of regression without outliers is up slight: $\alpha=0.754,0.772,0.770$, and $C$ is down to around 2.
3.2 The protein domains that are mostly responsible for the redundancy
Apparently, more tokens than types means that some CRCPDs appear multiple times in a chromosome. Which CRCPDs appear more than others? Table 1 shows the top three Pfam domains in each human chromosome, as well as in the whole genome, according to their frequency. On the whole genome level, zinc finger C2H2 type (zf-C2H2), a C2H2-type zinc finger (zf-C2H2_6), and Immunoglobulin I-set domain (I-set) are the most frequent domains, appearing 5493, 1609, and 900 times, respectively. Genomewide there are 6839 Pfam domain names, whose ID and annotation can be found in http://ftp.ebi.ac.uk/pub/databases/Pfam/current_release/Pfam-A.hmm.gz or http://pfam-legacy.xfam.org/family/browse .
On individual chromosome level, zf-C2H2 domain (Miller et al., 1985) is among the top 3 domains for 19 out of 23 chromosomes (not on the top 3 for chromosomes 2,11,14, and 22). Other top ranking domains on individual chromosome level include olduvai domain (chr1), immunoglobulin I-set domain (chr2), cadherin domain (chr4,5), 7 transmembrane receptor (rhodopsin family) (chr11), collagen triple helix repeat (chr13), immunoglobulin V-set domain (chr14, 22), and keratin high sulfur B2 protein (chr21).
3.3 Chromosome 19 as outlier
The chromosome 19 has already been known to have the highest gene density compared to other chromosomes, more gene families, high GC content, higher density of repetitive sequences (Grimwood et al., 2004), extreme divergence with mouse genome (Castresana, 2002) but conserved within nonhuman primates (Harris et al., 2020).
We examine several chromosome level statistics to see which quantities make chromosome 19 an outlier. The chr19 has the highest GC content compared to other chromosomes, which might be related to the higher gene density, which itself is a product of widespread DNA/gene duplication.
If CRCPD statistics are examined directly, we find that the proportion of unique CRCPD (number of CRCPD types whose token, or copy number, is equal to 1) in chr19 does not stand out (see Table 1). The number of CRCPD types on chr19 is large, but not the largest. However, chr19 has the highest average copy number per CRCPD type, and also has the largest copy number for a CRCPD: zinc finger C2H2 domain appears 2805 times, which is an order of magnitude larger than the maximum copy number in the other chromosomes.
3.4 Extension to half of the chromosome arms
The x-axis of Fig.1 only spans one decade (ratio of the maximum and minimum token count is 12 for chromosomes and 17 for chromosome arms). In order to increase the range of x-axis, we split each chromosome arm into two equally sized regions. The type-token with this relatively local data being added is shown in Fig.2.
Figure 2 shows that with more data at the low end of the x-axis (the largest token count is 59 times the smallest token count), Heaps’ law remains to be true, with the fitting exponent $\alpha=0.75$. Removing the outliers due to chr.19, $\alpha=0.8$. Both of these slopes are larger than those in Fig.1.
Visual inspection of Fig.2 seems to show a slight tendency for the points to curve down, which might explain the increase of the fitting exponent value. In fact, in the original version of Heaps’ law in Eq.(1), $C$ should be 1 as the line should go through $x=y=1$ points (when there is only one token, it must be unique). All our fitting $C$ values are larger than 1, meaning that in the log-log scale, type-token scatter plot will eventually deviate from a straight line towards the origin. We will return to this issue in section 5.
4 Results for Other Animal Genomes
Using the same UCSC Genome Browser, we downloaded the Pfam domain locations for mouse genome (GRCm39/mm39, Jun 2020, all 21 chromosomes), UniProt (The UniProt Consortium, 2012) domain information for chicken (GRCg6a/galGal6, 2018, missing chromosomes 29, 34-39), zebrafish (GRCz11/danRer11, 2017, all 25 chromosomes), and drosophila melanogaster (BDGP release 6, 2014, 2L, 2R, 3L, 3R, 4, X and Y chromosomes/arms).
Figure 3 shows the type-token scatter plot in log-log scale with points both representing a chromosome (black) or half of a chromosome (red). The practice of using half chromosomes is to increase the sample size and to increase the covering range of the x-axis.
The mouse genome exhibits similar behavior as the human genome. The variance explained by the power-law is $R^{2}=0.887$, the regression slope is $\alpha=0.688$ and $C=3.8$. The zebrafish genome has two chromosome outliers, chromosome 4 and 22, as well as half chromosome from chromosome 4. Removing these four points lead to a power-law regression with $R^{2}=0.816$ and slope=0.643, $C=3.62$. The chicken genome has two outliers at the chromosome level: chr16 and chr31, and 4 half-chromosome level outliers derived from these two chromosomes. After removing these six points, the regression in log-log scale has $R^{2}=0.937$, slope=0.782, and $C=1.56$. Finally, for drosophila genome, $R^{2}=0.993$, slope=0.865, and $C=0.81$.
To summarize, the results from these genomes where the protein domain track is available, Heaps’ law is generally true with exponent $\alpha<1$, despite potentially data quality issues such as the UniProt domain information, limited range of x-axis, and limited number of sample points. We also notice that when the low-end of the x-axis for the scatter plot is close to the origin (e.g. drosophila), the slope is relatively larger and closer to 1.
5 Artificial units generated by sampling tokens randomly and a systematic deviation from the power-law
The scatter plots in Fig.1 and Fig.2 are still too noisy, and the range of x-axis is still too limited. In order to have a clearer picture of the type-token relation in an even wider range, we cannot rely on the naturally provided units (“texts”) in chromosomes and chromosome arms, but to create artificial ones. We pool the appearance of CRCPD from all human chromosomes (62299 tokens). Then we randomly sample x number of tokens and check how many of them are unique (y number of types). A type-token plot from this random sampling is shown in Fig.4(A) in log-log scale.
If Heaps’ law holds exactly true, we should expect a straight line in the log-log plot in Fig.4. The linear regression (grey line) captures this trend (exponent 0.725 is less than one). However, a systematic deviation from the straight line, of a concave shape at both ends of the line, can be easily seen. When we add a quadratic term, it fits the data perfectly (with the coefficient for the quadratic term to be negative). The AIC of the quadratic regression, $-244.79$, is much better than that for the linear regression ($-93.08$), confirming the visual impression. This “log-log concavity” was previously reported (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015) and our DNA words result fully conforms it. For this reason, a better term for plots in Figs.1,2,3,4(A) could be “range-limited Heaps’ law”.
Since quadratic curves fit the type-token plot in Fig.4(A) nicely, the coefficient of the quadratic term is a reliable measure of the concavity. We investigate the impact of limited dictionary size on the concavity, or deviation from the Heaps’ law, by removing rare CRCPD. Fig.4(B) shows the fitting coefficient of the quadratic term at these situations: half of the singleton CRCPD types (those that appear in the genome only once) are removed; all singleton CRCPD types are removed; all singleton plus half of the doublet (those CRCPD types that appear in the genome only twice) are removed; and all singletons and all doublets are removed.
Each point in Fig.4(B) represents a random sampling process similar to that in Fig.4(A), i.e., a sequence of token sampling followed by type counting, when the type source (dictionary) is cut short. We ran the random sampling multiple times even with the same dictionary size, because of the variation of random sampling, as well as changing the upper limit of the x-axis to accommodate the reduction of dictionary size. Even with these variations, the trend is clear that the type-token plot becomes more curved if more rare words (singleton and doublet protein domains) are removed.
The random sampling of tokens can also be carried out on per chromosome basis. Fig.5 shows the type-token plot for 23 individual chromosomes. Protein domains observed only within one chromosome are pooled, then randomly sampled without replacement, and distinct types are counted. Chromosome 19 again stands out as the chromosome with the lowest diversity and lesser number of types. The plateaus reached for each chromosome indicate that the dictionary size has been reached. This finite size effect is different from the slight convexity of the type-token curve.
The lines/curves for individual chromosomes in Fig.5 behave similarly in terms of slope before reaching the plateau. The y-intercept is very different for chromosome 19, consistent with Fig.1 and Fig.2, while those of other chromosomes are comparable. The shape of per-chromosome type-token scatter plots obtained by random sampling, which is roughly a power-law with a slight concave curving followed by a plateau, is a typical example of finite size effect. Due to the nature of generated data, the number of token is not limited whereas the number of types is finite, leading to plateaus.
6 Discussion
Heaps (or Herdan-Heaps) law is one of the classic linguistic “laws” (Altmann and Gerlach, 2016; Hernández-Fernández et al., 2019). The word “law” in this context is very different from that in other fields such as physics (e.g. Newton’s laws). Linguistic laws are empirical, not supposed to be exact, and only describe a trend. As reported by several authors and confirmed in our own DNA word data, Heaps’ law or power-law relationship (or linear relation in log-log scale) is unlikely to be true for all range of text lengths (number of tokens) (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015). Even for human languages (e.g., some non-Indo-European languages), #type tends to flatten out away from the power-law in the large # token end (Lü et al., 2013). At low #token end, the fact that the fitting parameter C $>$ 1 (or C’ $>$ 0), in our data, indicates that the power-law will not be extrapolated all the way to the origin.
Since the genome size is limited, the total number of protein-domain-coding loci is limited, there is an upper limit in the #token axis. Similarly, it is not practical to use short genome regions to count #token and #type as half of the human genome are in non-coding intergenic regions. This leads to a lower limit in #token axis. Checking type-token relation in the human genome has to be range limited. On the other hand, our simulation in Fig.4 shows that beyond certain range, one should not expect power-law relationship between type and token anymore, and Heaps’ law will break down.
There are several publications relating exponents in Zipf’s law and Heaps’ law asymptotically if both the token frequency vs. rank and type vs. token are indeed inverse power-law and power-law (Baeza-Yates and Navarro, 2000). There are also papers deriving the asymptotic type-token relation in the model of random sampling from a power-law distributed word pool (Eliazar, 2011; Gerlach and Altmann, 2013; Boytsov, 2017), in complicated mathematical expressions often in the form of infinite summation. With finite dictionary size, infinite summation becomes finite summation, but still no simple mathematical expression is proposed. Here, from a data fitting perspective, the type-token relation can be expressed by a quadratic equation: $\log(y)=c^{\prime}+\alpha\log(x)-\beta(\log(x))^{2}$ with $10<\alpha<$ 1 and $\beta>0$. Using an extra term and extra parameter (Li et al., 2010; Li and Miramontes, 2011), in particular, using a quadratic/parabolic term (Frappat et al., 2003), to correct deviation from a straight line is a common practice.
Studies along the same lines with the present investigation include (Nasir et al., 2017; Caetano-Anollés, 2021). These works focus on phylogenetics of different species, and the word used is protein fold superfamilies (FSF). FSF information was provided by the SCOP (structural classification of proteins) database (Murzin et al., 1995; Andreeva et al., 2020). Even if SCOP domain superfamilies and Pfam domains are somewhat similar, the sample points in a type-token plot in (Nasir et al., 2017; Caetano-Anollés, 2021) are different from those in the current study. In (Nasir et al., 2017; Caetano-Anollés, 2021), each point represents a species whereas in our study a point is a chromosome or chromosome arm. When all species are represented in one plot in (Nasir et al., 2017; Caetano-Anollés, 2021), including viruses, archaea, bacteria, eukarya, the scattering of the points (genomes) in log-log scale does not follow one single straight line, but a curve.
We have confirmed the result of some previous works that if tokens are randomly sampled from a dictionary (type pool), while type frequency (token per type) follows the Zifp’s law or some modification of the Zipf’s law, the resulting type-token plots, though may have a look of power-law/Heaps’ law, are actually curved (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015). By simulation, we also numerically determined the impact of finite dictionary size, by removing the rare CRCPD types from consideration: the type-token plot becomes more curved with a reduction of dictionary size. The previous attempts to derive the type-token relation, with finite or infinite dictionary size do not present the result in closed-form analytic expression (Eliazar, 2011; Lü et al., 2013; Tunnicliffe and Hunter, 2022). Our quadratic regression provides a simple formula for an empirical description of the systematic deviation from the Heaps’ law. Adding a second term to better fit the result is a common practice in data analysis (Frappat et al., 2003; Li et al., 2010).
The definition of “type” or “dictionary word”, or in our case, CRCPD, is supposed to be given by the database, and synonym is not merged. For example, there are 13 protein domain names containing the prefix of zf-C2H2, including the top ranking zf-C2H2 (5493 appearances in the genome) and second ranking zf-C2H2_6 (1609 appearance in the genome). The reason of not combining potential synonyms is that it should be a feature of a “dictionary”, and actually, Heaps’ law partially reflects this lexical diversity. If one would combine CRCPD that are “similar” (up to certain degree), the type-token plot and Heaps’ law in human languages should also be modified accordingly.
As argued in (Lynch and Conery, 2003), a consequence of smaller population size is that selection pressure is much less effective in removing individuals without a perfect fit for the given environment. As nonadaptive forces dominated, genomes became less compact, with more redundancies, and inefficiencies were tolerated. On the other hand, this near-neutral random walk in Wright’s landscape allows a species to reach a new adaptive peak, potentially higher than the previous one (Wagner, 2019). Microorganisms have populations several orders of magnitude larger than animals. The higher level of genome redundancy and complexity in animals, versus the lower level of genome complexity in microorganisms, is then be expected. For a genome without redundancy, the number of token is roughly equal to the number of types, and Heaps exponent is one. With redundancy, exponent is expected to be less than 1.
It is suggested in (Caetano-Anollés et al., 2017; Caetano-Anollés, 2021) that viral, archaeal and bacterial proteomics have a higher level of vocabulary growth than eukarya, as indicated by the higher value of $\alpha$ exponents (around 0.8). The $\alpha$ exponent for eukarya genomes in (Caetano-Anollés, 2021) is only around 0.1-0.3, in contrast with our exponent $\alpha\sim 0.7$ within one genome. Because the higher the amount of duplication and larger the resulting gene families, the less the type-to-token ratio, we expect genomes like the human genome with abundance of duplication to have a lower $\alpha$ exponent than genomes lacking duplication. In our data and result, human, mouse, and zebrafish exhibit a lower value of $\alpha$ than those for chicken and drosophila. This lower $\alpha$ exponent is indicative of a larger amount of duplicated protein domains which is typical in presumably higher organisms.
[…]
Despite the deviation from a perfect power-law for type-token plot, or in an alternative wording, that the power-law Heaps’ law is only technically correct in a limited range, there are other important lessons in this study from the linguistic perspective. Just like human languages with hundreds of new English words being created annually, albeit some of them are synonyms to the existing words, new protein domains have been generated during evolution. Even if the new protein domains may not be completely new, being at various degress similar to the old ones, these may not be that different from the new synonymous words in human languages. Some of the seemingly non-novel new protein domains may continue to evolve and eventually lead to a new meaning and a new structure.
Acknowledgment
We would like to thank Patrick Villanueva for participating to the initial stage of the project, Tom MacCarthy and Roman Samulyak for useful suggestions.