### v2

# Range-Limited Heaps’ Law for Functional DNA Words in the Human Genome

Abstract
Heaps’ or Herdan’s law is a linguistic law describing the relationship between the vocabulary/dictionary size (type) and word counts (token) to be a power-law function. Its existence in genomes with certain definition of DNA words is unclear partly because the dictionary size in genome could be much smaller than that in a human language. We define a DNA word as a coding region in a genome that codes for a protein domain. Using human chromosomes and chromosome arms as individual samples, we establish the existence of Heaps’ law in the human genome within limited range. Our definition of words in a genomic or proteomic context is different from other definitions such as over-represented k-mers which are much shorter in length. Although an approximate power-law distribution of protein domain sizes due to gene duplication and the related Zipf’s law is well known, their translation to the Heaps’ law in DNA words is not automatic. Several other animal genomes are shown herein also to exhibit range-limited Heaps’ law with our definition of DNA words, though with various exponents. When tokens were randomly sampled and sample sizes reach to the maximum level, a deviation from the Heaps’ law was observed, but a quadratic regression in log-log type-token plot fits the data perfectly. Investigation of type-token plot and its regression coefficients could provide an alternative narrative of reusage and redundancy of protein domains as well as creation of new protein domains from a linguistic perspective.
Abbreviations: AIC: Akaike information criterion; CRCPD: coding regions that code for protein domains; GC content: percentage of guanine and cytosine bases.
1 Introduction
The study of molecular biology is full of examples in terminology, and in quantitative and empirical laws, borrowed from linguistics (Key, 2000). Jargons used in molecular biology, such as transcription, translation, code, etc. are often borrowed from terms in linguistics, not to mention the ubiquitous use of letters (nucleotides and amino acids) and text (DNA and protein sequences). The recent Nobel-prize-winning biotechnique of CRISPR was described as an equivalence of a text editor (Nelson et al., 2015). Not surprisingly, linguistic laws have also seen their mentioning in genomics. For example, Zipf’s law (Zipf, 1935) which describes how the word frequency in a text drops off with its rank, was tested in DNA sequences with k-mers/k-tuples (e.g. k=6) considered as “word” (Mantegna et al., 1994). This study was criticized, putting aside the issue of the definition of DNA words, because the rank-frequency plot did not show a good power-law as in the case of human languages (Konopka and Martindale, 1995; Li, 2002). Another example of linguistic law is the Menzerath’s law (Menzerath, 1928), which describes the tendency for using smaller linguistic units when these units are more numerous. The Menzerath’s law at the human exon level is the tendency for genes with more exons to have smaller exon sizes (Li, 2011a; Nikolaou, 2014), while other choices of units are also investigated (Semple et al., 2022).
In this study, we focus on another linguistic law, the Heaps law (Heaps, 1978) (or Herdan-Heaps law (Egghe, 2007)), which states that the number of unique words (vocabulary/dictionary size) as a function of text length is a power-law. Heaps’ law can be understood as a (number of)type-(number of)token (Herdan, 1960; Wetzel, 2009) relationship: type being a unique word, token being usage of a word, i.e., the total number of types is the vocabulary size, and total number of tokens is the vocabulary usage. Heaps’ law can also be understood as a law of diminishing return (Petersen et al., 2012). If one searches longer and longer texts for the purpose of finding new words, Heaps’ law indicates that the chance of finding new words becomes smaller and smaller as more texts are looked over. The latter meaning in Heaps law has been implied in the number of genes or gene-families discovered by sequencing more and more species’ genomes in a clade (Medini et al., 2020), or the number of single-nucleotide-polymorphisms (SNPs) discovered by sequencing more samples (Ionita-Laza et al., 2009).
In order to apply a linguistic law to genomics, one first needs to find an entity corresponding to a word. A common practice is to consider (overlapping) k-mers with a prefixed k value as “words” due to its easiness to calculate, its use in alignment-free sequence comparison (Zielezinski et al., 2017; Li et al., 2019), and its importance in many bioinformatics algorithms (Chikhi and Medvedev, 2013; Koren et al., 2017; Rahman et al., 2018). However, the problem with all fixed-length k-mers as words is that k-mers in general lack functional meaning.
Instead of all k-mers, attempts were made to select only the over-represented k-mers, based on the frequency expectation from (k-1)-mers, as words or motifs (Brendel et al., 1986; Phillips, et al., 1987; Vilo, 2002; Apostolico et al., 2003; Gatherer, 2007). There are other motif detection methods based on statistical mechanics (Bussemaker et al., 2000; Moghaddasi et al., 2017). However, these are postulated units only. In fact, non-coding regions tend to have relatively more over-represented oligonucleotides than coding regions (Frontali and Pizzi, 1999). Many of over-represented oligonucleotides in noncoding regions may be due to error accumulation (e.g. slippage errors) tolerated by purifying selection, but some others may have functional implications and thus deserve the name “words”, as the mirror-symmetrical words found in introns (Brendel et al., 1986).
In the framework of semiotics (Hoopes, 2011), word is part of the signifier-signified pair that represents a meaning. Without a biological meaning, a proposed DNA or protein word would be much less interesting if we continue to discuss linguistic laws. Although the in-frame non-overlapping 3-mers within exons (codons) do have a biological meaning (Mukhopadhyay et al., 2006), k-mers, even the over-represented ones, may or may not.
Recently, large language models (Brant et al., 2007), such as bidirectional encoder representations from transformers (BERT) (Devlin et al., 2018) or generative pre-trained transformer (GPT) (Radford et al., 2019), have been applied to biomolecular sequences, in particular protein sequences (Rao et al., 2020; Madani et al., 2023; Nijkamp et al., 2023). One aspect of natural language process is tokenization which means to delineate a character string by boundaries (Webster and Kit, 1992). For protein sequences, the tokenization is a process to find “words” (Ofer et al., 2021). The protein “words” found by this method are usually short, with average size of only a few (e.g. four) amino acids (Ferruz et al., 2022; Dotan et al., 2024).
In this study, we use DNA segments that code a protein domain (Wang et al., 2021) as our DNA “word tokens”. Treating protein domains as protein words has been proposed before (Searls, 2002; Gimona, 2006; Scaiewicz and Levitt, 2015; Yu et al., 2019; Buchan and Jones, 2020). There were studies that consider gene internal structure in DNA sequences as “grammar” or rules (Dong and Searls, 1994), but the emphasis of these were not on the definition of words. We prefer to define DNA words as substrings of gene sequences delimited by the boundaries of the protein domains within the proteins encoded by those genes. By doing so, our definition of DNA words has a biological meaning, i.e. coding for a protein domain.
To investigate if Heaps’ law holds in the human genome, our task is to find all Coding Regions that Code for a Protein Domain (CRCPD). For this we need to, for any genomic region, count the number of all CRCPDs (number of tokens) and the number of distinct CRCPDs among them (number of types). For one chromosome of the human genome, this pair of counts (#token, #type) will only give us one point. To see a relationship between token and type, we repeat the same counting process for each one of the 23 chromosomes and/or each one of the chromosome arms, so more points are produced. This may be chosen if more points are desired, as chromosomal arms may be seen as relatively independent entities, and different from the whole chromosomes where they belong.
It has been known for over twenty years that the histogram of protein domains or distinct parts follow an inverse power-law distribution in genomes (Qian et al., 2001; Luscombe et al., 2002; Koonin et al., 2002). Another well known property is the importance role that gene duplication plays in shaping the distribution of these protein domains as well as other genomic units (Müller et al., 2002; Gao and Miller, 2011; Li et al., 2016), echoing theoretical studies linking duplication and power-law behavior (Li, 1991; Ispolatov et al., 2005). A perfect inverse power-law histogram (distribution) will lead to a similar inverse power-law with rank as the x-variable, i.e., the Zipf’s law (Newman, 2005).
However, more and more studies show that when tokens are randomly chosen from a dictionary with a finite size whose words follow Zipf’s law distribution, the resulting type-token plot does not follow an exact Heaps’ law (Bernhardsson et al., 2009; Font-Clos and Corral, 2015). In particular, the type-token plot in log-log scale would show a concave downward shape (Font-Clos and Corral, 2015). Nevertheless, within a limited range, this concave curve might be approximately viewed as a straight line. Also, publications with proofs on relations between two power-law exponents in the two laws rely on certain assumptions, e.g. infinite dictionary size, and working in the infinite token limit (Boytsov, 2017). All these observations imply that the power-law distribution in protein domain known for the last twenty years does not automatically lead to a power-law type-token plot for CRCPD, which needs to be checked from the data directly.
In the next section, we describe the data and the methodology used. In section 3, we present evidence of the existence of range-limited Heaps’ law in the human genome; while we point out one chromosome (chromosome 19) as an outlier to the Heaps’ law; We also include in our study another level of genomic entities, chromosome arms, to explore the extent of Heaps’ law. In section 4, we examine the potential Heaps’ law in other animal genomes. In section 5, we generate artificial units by randomly sampling DNA word tokens from the list of known DNA word types, and confirm the previous results of log-log concave downward and further reconfirm a systematic deviation from the Heaps’ law. The same random sampling process is also carried out at the individual chromosome level, showing a distinction between the concavity and finite dictionary size effect. Finally, various points are discussed in the Discussion section, including a comparison between the creation of new words in human languages and evolution of protein domains in genomes.
2 Data and Methods
The protein domains used in the present study refer to Pfam (Mistry et al., 2021) (https://pfam.xfam.org/ or https://pfam-legacy.xfam.org/, which is hosted by InterPro (Paysam-Lafosse et al., 2022) after January 2023: https://www.ebi.ac.uk/interpro). The chromosomal locations of CRCPD in human genes are available from the University of California at Santa Cruz (UCSC) Genome Browser track: Pfam domains in GENECODE genes. A description of the methods used to derive this track can be found on UCSC Genome Browser site (https://genome.ucsc.edu/cgi-bin/hgTrackUi?db=hg19&g=ucscGenePfam).
To obtain this Pfam domains track in the human genome (for GRCh38/hg38), from the top bar of the genome browser, we select Tools $\rightarrow$ Table Browser, group = “Genes and Gene Predictions”, track = “Pfam in GENECODE”. The resulting table contains chromosome number, chromosomal location (start, end, in bp), name of the Pfam domain, etc.
In this study, we also used chromosome arms as units. To determine the chromosome arm to which a region belongs to, we download the cytoband information from:
https://hgdownload.cse.ucsc.edu/goldenpath/hg38/database/cytoBand.txt.gz to find the centromere locations in each chromosome. The counting of Pfam domains (both token and type) was carried out by R (https://www.r-project.org) scripts written by us.
The Heaps’ law is written as ($y$= vocabulary size = #type, or the number of unique CRCPD domains in a chromosome, $x$=vocabulary usage=#token, or number of times any of these CRCPD appear in the same chromosome):
$y=Cx^{\alpha}$ (1)
which after a logarithmic transformation on both side ($y^{\prime}=log(y)$, $x^{\prime}=log(x)$, $C^{\prime}=log(C)$): $y^{\prime}=C^{\prime}+\alpha x^{\prime}$ becomes a linear relationship. The exponent $\alpha$ is determined by a linear regression in R (the function lm). For model selection the Akaike information criterion (AIC) (Akaike, 1974; Li and Nyholt, 2001) calculation is carried out by the AICcmodavg package (cran.r-project.org/web/packages/AICcmodavg/), using the function aictab. For quadratic regression, we fit the equation $y^{\prime}=C+\alpha x^{\prime}+\beta x^{\prime 2}$.
The following notation may gives readers an idea of the type of data we have. Suppose there are N tokens (appearance of CRCPDs: coding for Pfam domains), and K unique ones. $N=\sum_{i=1}^{K}P_{k}$, where $P_{k}$ is the multiplicity of word type-$k$. For any genomic region (chromosome, chromosome arm, or half-arm), we count the number of tokens $x_{i}$ ($i$ is the index for sample point) and $y_{i}$ of them are unique. When we have enough samplings $i=1,2,\cdots,I$, these ($x_{i}$, $y_{i}$) pairs can be log-transformed and used for a linear regression. We may not be constrained by genomic regions, but sample $x_{i}$ tokens genome-wide from the pool of N tokens (sampling without replacement) and among them $y_{i}$ ones are unique. This can be repeated I times with ever larger $x_{i}$ values. The same ($x_{i}$,$y_{i}$) pairs can be used to check Heaps’ law.
3 Results for the human genome
3.1 Heaps’ law for Pfam CRCPD holds true
The number of Pfam domain appearance, as our example of CRCPD, in each autosomal and X chromosome in the human genome (version hg38) is obtained from the Pfam-in-GENECODE track of the UCSC genome browser. This would produce 23 points for checking the vocabulary size (type) and vocabulary usage (token) relationship and a regression. In order to increase the sample size, we also consider data from the p- and q-arm in a chromosome, as partitioned by the centromere for each metacentric (non-acrocentric) chromosome (41 points).
Fig.1(A) shows the number of types (number of unique/distinct Pfam domains) as a function of number of tokens (Pfam domain counts) for 23 human chromosomes in log-log scale. With the exception of one point, all other points scatter around a straight-line, indicating a power-law and Heaps’ law. The outlier is chromosome 19, being lower in number of CRCPD types or higher in number of CRCPD tokens. The regression for log-token over log-type produces the slope (scaling exponent) $\alpha=0.658$ and $C=5.12$ in Eq.(1).
Fig.1(B) shows a similar type-token plot in log-log scale using chromosome arm data. The regression results are similar: $\alpha=0.683$ and $C=3.97$. Using both whole chromosome and chromosome arm data leads to Fig.1(C), with $\alpha=0.695$ and $C=3.7$. Figure 1(D) shows the same data as Fig.1(C) but in linear-linear scale. It is clear that power-law curve function fits the data better than a linear regression. In fact, the variance explained by regression is much higher in log-log scale ($R^{2}=0.887,0.873,0.9$ for chromosome data, chromosome arm data, and combined) than in linear-linear scale ($R^{2}=0.697,0.667,0.738$).
Figure 1: Type-token plot, where $x$-axis is the number of CRCPD tokens (vocabulary usage) and $y$-axis is the number of unique CRCPD type (vocabulary size), for human genome. (A) Each point corresponds to a human chromosome, in log-log scale. (B) Each point corresponds to a human chromosome arm, in log-log scale. (C) Each point corresponds to either a human chromosome or a chromosome arm, in log-log scale. (D) Same as (C) but in linear-linear scale. Chromosome 19 is clearly indicated as an outlier in all panels.
Removing the outliers (one point in Fig.1(A), two points in Fig.1(B) for two chromosome 19 arms, and three points in Fig. 1(C)) result in even better $R^{2}$: 0.968, 0.955, and 0.967. The slope of regression without outliers is up slight: $\alpha=0.754,0.772,0.770$, and $C$ is down to around 2.
3.2 The protein domains that are mostly responsible for the redundancy
Apparently, more tokens than types means that some CRCPDs appear multiple times in a chromosome. Which CRCPDs appear more than others? Table 1 shows the top three Pfam domains in each human chromosome, as well as in the whole genome, according to their frequency. On the whole genome level, zinc finger C2H2 type (zf-C2H2), a C2H2-type zinc finger (zf-C2H2_6), and Immunoglobulin I-set domain (I-set) are the most frequent domains, appearing 5493, 1609, and 900 times, respectively. Genomewide there are 6839 Pfam domain names, whose ID and annotation can be found in http://ftp.ebi.ac.uk/pub/databases/Pfam/current_release/Pfam-A.hmm.gz or http://pfam-legacy.xfam.org/family/browse .
On individual chromosome level, zf-C2H2 domain (Miller et al., 1985) is among the top 3 domains for 19 out of 23 chromosomes (not on the top 3 for chromosomes 2,11,14, and 22). Other top ranking domains on individual chromosome level include olduvai domain (Chr1), immunoglobulin I-set domain (Chr2), cadherin domain (Chr4,5), 7 transmembrane receptor (rhodopsin family) (Chr11), collagen triple helix repeat (Chr13), immunoglobulin V-set domain (Chr14, 22), and keratin high sulfur B2 protein (Chr21).
ch most_freq CN 2nd CN 3rd CN num_type per(CN=1) mean(CN) max(CN)
1 Olduvai 215 zf-C2H2 214 Sushi 186 1759 64.7% 3.597 215
2 I-set 224 fn3 156 Ig_3 136 1355 63.8% 3.373 224
3 zf-C2H2 217 WD40 72 Ig_3 70 1178 65.3% 2.93 217
4 Cadherin 101 zf-C2H2 92 Ank 50 857 67.1% 2.797 101
5 Cadherin 334 zf-C2H2 134 Cadherin_2 55 951 68.2% 3.121 334
6 zf-C2H2 162 fn3 62 zf-C2H2_6 57 1057 65.8% 2.807 162
7 zf-C2H2 324 zf-C2H2_6 91 V-set 58 957 66.1% 2.883 324
8 zf-C2H2 171 Sushi 56 Ank 52 787 67.9% 2.663 171
9 zf-C2H2 144 I-set 57 Ig_3 57 856 67.4% 2.974 144
10 zf-C2H2 103 Ank 72 Ank_3 44 854 66% 2.56 103
11 7tm_1 196 7tm_4 180 I-set 70 1136 65.8% 3.109 196
12 zf-C2H2 127 WD40 48 Ank 40 1100 65.8% 2.576 127
13 Collagen 28 Cadherin 21 zf-C2H2 19 475 69.7% 2.027 28
14 V-set 96 WD40 41 7tm_1 40 736 67.1% 2.455 96
15 zf-C2H2 57 EGF_CA 46 hEGF 41 727 68.1% 2.611 57
16 zf-C2H2 292 zf-C2H2_6 71 WD40 44 914 67.2% 2.534 292
17 zf-C2H2 107 Keratin_B2_2 73 WD40 42 1151 65.2% 2.487 107
18 zf-C2H2 75 Cadherin 45 Laminin_EGF 30 370 69.5% 2.473 75
19 zf-C2H2 2805 zf-C2H2_6 910 KRAB 228 1114 65.4% 6.684 2805
20 zf-C2H2 107 zf-C2H2_6 25 Laminin_EGF 21 674 72.6% 2.068 107
21 Keratin_B2_2 47 I-set 17 ig 16 291 69.4% 2.131 47
22 V-set 39 zf-C2H2 27 EGF_CA 16 590 68.1% 2.171 39
X zf-C2H2 136 Collagen 32 MAGE 32 781 59.3% 2.77 136
all zf-C2H2 5493 zf-C2H2_6 1609 I-set 900 6839 44.3% 9.109 5483
Table 1: Some human chromosome level statistics on protein domains (CN: copy number, or number of CRCPD token for a specific CRCPD type): the protein domain with the highest, second highest, and the third highest, CN; number of unique CRCPDs; proportion of CRCPD types that have CN=1 (which could be called a singleton); average copy number (or token-over-type ratio); and maximum copy number. The maximum (minimum) value in each column is marked with bold (italics).
3.3 Chromosome 19 as outlier
Chromosome 19 (Chr19) has already been known to have the highest gene density compared to other chromosomes, more gene families, high GC content, higher density of repetitive sequences (Grimwood et al., 2004), extreme divergence with mouse genome (Castresana, 2002) but conserved within nonhuman primates (Harris et al., 2020).
We examine several chromosome level statistics to see which quantities make chromosome 19 an outlier. Chr19 has the highest GC content compared to other chromosomes, which might be related to the higher gene density, which itself is a product of widespread DNA/gene duplication.
If CRCPD statistics are examined directly, we find that the proportion of unique CRCPD in Chr19 does not stand out (see Table 1). The number of CRCPD types on Chr19 is large, but not the largest. However, Chr19 has the highest average copy number per CRCPD, and also has the largest copy number for a CRCPD: zinc finger C2H2 domain appears 2805 times, which is an order of magnitude larger than the maximum copy number in the other chromosomes.
3.4 Extension to half of the chromosome arms
The x-axis of Fig.1 only spans one decade (ratio of the maximum and minimum token count is 12 for chromosomes and 17 for chromosome arms). In order to increase the range of x-axis, we split each chromosome arm into two equally sized regions. The type-token with this relatively local data being added is shown in Fig.2.
Figure 2 shows that with more data at the low end of the x-axis (the largest token count is 59 times the smallest token count), Heaps’ law remains to be true, with the fitting exponent $\alpha=0.75$. Removing the outliers due to Chr19, $\alpha=0.8$. Both of these slopes are larger than those in Fig.1.
Visual inspection of Fig.2 seems to show a slight tendency for the points to curve down, which might explain the increase of the fitting exponent value. In fact, in the original version of Heaps’ law in Eq.(1), $C$ should be 1 as the line should go through $x=y=1$ points (when there is only one token, it must be unique). All our fitting $C$ values are larger than 1, meaning that in the log-log scale, type-token scatter plot will eventually deviate from a straight line towards the origin. We will return to this issue in section 5.
Figure 2: An extension of Fig.1(C). Besides human chromosomes and chromosome arms, half-chromosome-arms (blue) are used to count CRCPD tokens and types.
4 Robustness of the results
4.1 Random shuffle of CRCPDs
In order to check if the sampling units of chromosomes, chromosome arms, or half arms, play any role in the observed power-law trend in Figs.1,2, we carried out two experiments. The first is that we shuffle the CRCPDs (tokens) genome-wide once, while keeping the chromosome location intact. Then we repeat the same CRCPD-type count in the same chromosomes, chromosome arms, and half arms. In the next subsection, we will randomly sample CRCPDs without replacement. Then the CRCPD-type count is plotted against the number of CRCPD token sampled.
Fig.3 shows the type-token relationship when CRCPDs are shuffled in the whole genome. The black dots are identical to those in Fig.2, some dots represent chromosomes, others represent chromosome arms or half arms. The red dots are the similar type counts in the same genomic segments (chromosomes, arms, half arms) when the shuffling has taken place.
It can be seen from Fig.3 that after shuffling, each genomic unit tends to have more CRCPD types than the original data (red dots being above black dots). It means that CRCPD redundancy tends to be locally clustered. However, the shuffling of CRCPDs does not change the fact that type and token are in a power-law relationship. The regression coefficient is 0.747 for the shuffled data, very close to the original values of 0.747 (or 0.8 when Chr19 is removed) for the original data.
Figure 3: CRCPD types count in chromosomes, chromosome arms, and half arms after CRCPD tokens are shuffled genome-wide (red dots), as a function of CRCPD token counts in these genomic units. The black dots are idential to those in Fig.2, for the same type-token counts when CRCPDs are not shuffled.
4.2 Random sampling from CRCPD pool and a systematic deviation from the power-law
In the second approach to test the robustness of our result, we create artificial sampling units by randomly picking CRCPD tokens from the genome. This new way of sampling tokens would free us from the restriction of chromosomes or any genomic regions. Also, the range of x-axis in type-token plot can be greatly expanded. We pool CRCPDs from all human chromosomes (62299 tokens). Then we randomly sample x number of tokens and check how many of them are unique (y number of types). A type-token plot from this random sampling is shown in Fig.4(A) in log-log scale. Note that the x-range in Fig.4(A) is much more expanded compared to Figs.1,2.
If Heaps’ law holds exactly true, we should expect a straight line in the log-log plot in Fig.4. The linear regression (grey line) captures this trend (exponent 0.725 is less than one). However, a systematic deviation from the straight line, of a concave shape at both ends of the line, can be easily seen. When we add a quadratic term, it fits the data perfectly (with the coefficient for the quadratic term to be negative). The AIC of the quadratic regression, $-244.79$, is much better than that for the linear regression ($-93.08$), confirming the visual impression. This “log-log concavity” was previously reported (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015) and our DNA words result fully conforms it. For this reason, a better term for plots in Figs.1,2,6,4(A) could be “range-limited Heaps’ law”.
Since quadratic curves fit the type-token plot in Fig.4(A) nicely, the coefficient of the quadratic term is a reliable measure of the concavity. We investigate the impact of limited dictionary size on the concavity, or deviation from the Heaps’ law, by removing rare CRCPD. Fig.4(B) shows the fitting coefficient of the quadratic term at these situations: half of the singleton CRCPD types (those that appear in the genome only once) are removed; all singleton CRCPD types are removed; all singleton plus half of the doublet (those CRCPD types that appear in the genome only twice) are removed; and all singletons and all doublets are removed.
Each point in Fig.4(B) represents a random sampling process similar to that in Fig.4(A), i.e., a sequence of token sampling followed by type counting, when the type source (dictionary) is cut short. We ran the random sampling multiple times even with the same dictionary size, because of the variation of random sampling, as well as changing the upper limit of the x-axis to accommodate the reduction of dictionary size. Even with these variations, the trend is clear that the type-token plot becomes more curved if more rare words (singleton and doublet protein domains) are removed.
The random sampling of tokens can also be carried out on per chromosome basis. Fig.5 shows the type-token plot for 23 individual chromosomes. Protein domains observed only within one chromosome are pooled, then randomly sampled without replacement, and distinct types are counted. Chromosome 19 again stands out as the chromosome with the lowest diversity and lesser number of types. The plateaus reached for each chromosome indicate that the dictionary size has been reached. This finite size effect is different from the slight convexity of the type-token curve in Fig.4.
The lines/curves for individual chromosomes in Fig.5 behave similarly in terms of slope before reaching the plateau. The y-intercept is very different for chromosome 19, consistent with Fig.1 and Fig.2, while those of other chromosomes are comparable. The shape of per-chromosome type-token scatter plots obtained by random sampling, which is roughly a power-law with a slight concave curving followed by a plateau, is a typical example of finite size effect. Due to the nature of generated data, the number of tokens is not limited whereas the number of types is finite, leading to plateaus.
Figure 4: (A) Token(x-axis)-type(y-axis) plot for randomly generated token from the whole token pool of 62299 CRCPDs. The plot is fitted by both linear regression (after logarithm transformation of both axes) and quadratic regression. Quadratic regression visibly fits the data better than linear regression, indicating a deviation from the Heaps’ law. The data points from Fig.1(C) are redrawn here (lightblue) for comparison (for both chromosomes and chromosome arms, except for Chr19). (B) Repeating the same random sampling of token and quadratic regression fitting, when the dictionary size is gradually reduced. The y-axis shows the number of CRCPD type (from top to bottom): original data; half of the singleton CRCPD types (those that appear in the human genome only once) are removed; all singletons are removed; all singletons and half of the doublet CRCPD types are removed; all singletons and all doublets are removed. The x-axis is the regression coefficient for the quadratic term.
Figure 5: Type-token plots for randomly sampled CRCPDs in individual chromosomes. Note that Chr19 (lower purple curve) distinguishes itself from the rest (see discussion in the text).
5 Results for other animal genomes
5.1 Sampling from chromosomal units
Using the same UCSC Genome Browser, we downloaded the Pfam domain locations for mouse genome (GRCm39/mm39, Jun 2020, all 21 chromosomes), UniProt (The UniProt Consortium, 2012) domain information for chicken (GRCg6a/galGal6, 2018, missing chromosomes 29, 34-39), zebrafish (GRCz11/danRer11, 2017, all 25 chromosomes), and drosophila melanogaster (BDGP release 6, 2014, 2L, 2R, 3L, 3R, 4, X and Y chromosomes/arms).
Figure 6 shows the type-token scatter plot in log-log scale with points both representing a chromosome (black) or half of a chromosome (blue). The practice of using half chromosomes is to increase the sample size and to increase the covering range of the x-axis.
The mouse genome exhibits similar behavior as the human genome. The variance explained by the power-law is $R^{2}=0.887$, the regression slope is $\alpha=0.688$ and $C=3.8$. The zebrafish genome has two chromosome outliers, chromosome 4 and 22, as well as half chromosome from chromosome 4. Removing these four points lead to a power-law regression with $R^{2}=0.816$ and slope=0.643, $C=3.62$. The chicken genome has two outliers at the chromosome level: Chr16 and Chr31, and 4 half-chromosome level outliers derived from these two chromosomes. After removing these six points, the regression in log-log scale has $R^{2}=0.937$, slope=0.782, and $C=1.56$. Finally, for drosophila genome, $R^{2}=0.993$, slope=0.865, and $C=0.81$.
To summarize, the results from these genomes where the protein domain track is available, Heaps’ law is generally true with exponent $\alpha<1$, despite potentially data quality issues such as the UniProt domain information, limited range of x-axis, and limited number of sample points. We also notice that when the low-end of the x-axis for the scatter plot is close to the origin (e.g. drosophila), the slope is relatively larger and closer to 1. Another observation is the difference of vocabulary sizes among different genomes, with mouse having more protein domain types than chicken, zebrafish and drosophila. Word size differences between genomes were also observed in (Gatherer, 2007), though for different definition of words (over-represented protein substrings).
Figure 6: Token-type plots for mouse, zebrafish, chicken, and drosophila genomes. Each black dot represents a chromosome and blue dot for half-chromosome. The chromosome number of some outliers are indicated. The regression slopes from log-log plot are the Heaps’ law exponent.
5.2 Random sampling from total CRCPD pool
Similar to subsection 4.2, we may create our own sampling units by randomly sampling CRCPD tokens from a genome. This step is even more important for other animal genomes than the human genome, because the type-token plots in Fig.6 are more noisy and are more limited in x-axis range.
Figure 7 shows the number of CRCPD types within a randomly sampled CRCPD tokens (without replacement), as a function of number of tokens, for the four genomes (mouse, drosophila, chicken, and zebrafish). The trend of these type-token plots is very similar: quadratic regression lines fit the data perfectly, and within limited range, these quadratic lines can be approximated as linear segments (thus Heaps’ law).
The quadratic nature of the type-token plot when x-axis is extended reveals a weakness of using linear regression to fit the data in Fig.6. The slight difference between the linear coefficients in the different genomes in Fig.6 is not a reliable index to rank genomes. Figure 7, on the other hand, shows a clear order, from zebrafish, to chicken and drosophila, then to mouse, following the trend of increase in vocabulary sizes (number of distinct CRCPD types). The human data is also shown in Fig.7 which is essentially identical to that of mouse.
Figure 7: For the four animal genomes (mouse, drosophila, chicken, and zebrafish), number of unique CRCPD types as a function of the number of randomly sampled CRCPD tokens (in log-log scale). The quadratic regression lines (for log-x and log-y) are also shown.
6 Discussion
Heaps (or Herdan-Heaps) law is one of the classic linguistic “laws” (Altmann and Gerlach, 2016; Hernández-Fernández et al., 2019). The word “law” in this context is very different from that in other fields such as physics (e.g. Newton’s laws). Linguistic laws are empirical, not supposed to be exact, and only describe a trend. As reported by several authors and confirmed in our own DNA word data, Heaps’ law or power-law relationship (or linear relation in log-log scale) is unlikely to be true for all range of text lengths (number of tokens) (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015). Even for human languages (e.g., some non-Indo-European languages), #type tends to flatten out away from the power-law in the large # token end (Lü et al., 2013). At low #token end, the fact that the fitting parameter C $>$ 1 (or C’ $>$ 0), in our data, indicates that the power-law will not be extrapolated all the way to the origin.
Since the genome size is limited, the total number of protein-domain-coding loci is limited, therefore there is an upper limit in the #token axis as well as in the #type axis. Similarly, due to the large proportion ($\sim$ 50%) of human genome being in non-coding intergenic regions without any CRCPDs, it is not practical to use a short genome region as it may have zero count of #token and #type. This leads to a lower limit in #token axis. Checking type-token relation in the human genome has to be range limited. On the other hand, our simulation in Fig.4 shows that beyond certain range, one should not expect power-law relationship between type and token anymore, and Heaps’ law will break down.
There are several publications relating exponents in Zipf’s law and Heaps’ law asymptotically if both the token frequency vs. rank and type vs. token are indeed inverse power-law and power-law (Baeza-Yates and Navarro, 2000). There are also papers deriving the asymptotic type-token relation in the model of random sampling from a power-law distributed word pool (Eliazar, 2011; Gerlach and Altmann, 2013; Boytsov, 2017), in complicated mathematical expressions often in the form of infinite summation. With finite dictionary size, infinite summation becomes finite summation, but still no simple mathematical expression is proposed. Here, from a data fitting perspective, the type-token relation can be expressed by a quadratic equation: $\log(y)=c^{\prime}+\alpha\log(x)-\beta(\log(x))^{2}$ with $10<\alpha<$ 1 and $\beta>0$. Using an extra term and extra parameter (Li et al., 2010; Li and Miramontes, 2011), in particular, using a quadratic/parabolic term (Frappat et al., 2003), to correct deviation from a straight line is a common practice.
Studies along the same lines with the present investigation include (Nasir et al., 2017; Caetano-Anollés, 2021). These works focus on phylogenetics of different species, and the word used is protein fold superfamilies (FSF). FSF information was provided by the SCOP (structural classification of proteins) database (Murzin et al., 1995; Andreeva et al., 2020). Even if SCOP domain superfamilies and Pfam domains are somewhat similar, the sample points in a type-token plot in (Nasir et al., 2017; Caetano-Anollés, 2021) are different from those in the current study. In (Nasir et al., 2017; Caetano-Anollés, 2021), each point represents a species whereas in our study a point is a chromosome or chromosome arm or a unit smaller than the whole genome. When all species are represented in one plot in (Nasir et al., 2017; Caetano-Anollés, 2021), including viruses, archaea, bacteria, eukaryotes, the scattering of the points (genomes) in log-log scale does not follow one single straight line, but a curve. Instead of treating the curve as consisting of multiple piecewise linear regimes (Nasir et al., 2017; Caetano-Anollés, 2021), another possibility is to use a quadratic fitting function; admittedly, this does not provide an explanation for the reason of deviation from the power-law trend, and it ignores some fundamental differences in genome evolution in different domains.
One key lesson from the Heaps’ law that the rate of new knowledge gain does not linearly increase with the effort, is very relevant to pan-genome studies (Tellelin et al., 2008) – the number of new genes discovered does not increase linearly with the number of genome sequenced. A different kind of diminishing return, concerning the alignment of next-generation-sequencing reads when the read length is increased, is discussed in (Li et al., 2014; Stephens and Iyre, 2018).
Besides the above work on protein fold superfamilies in pan-genomes, Heaps’ law, unlike Zipf’s law (Sheinman et al., 2016), rarely appears in the literature concerning genomics. In (Gatherer, 2007), when over-represented peptide substrings (over-represented k-mers with k in a range of values) are considered as words, linear regression is used to fit the type-token scatter plots. No attempt was made to use the power-law function in (Gatherer, 2007). In (Mukhopadhyay et al., 2006), the fixed-length (k=3) codons are considered as words. This definition of words would easily reach the vocabulary limit when the text length is increased. Therefore, it is not surprising that a different type-token relation is expected.
We have confirmed the result of some previous works that if tokens are randomly sampled from a dictionary (type pool), while type frequency (token per type) follows the Zifp’s law or some modification of the Zipf’s law, the resulting type-token plots, though may have a look of power-law/Heaps’ law, are actually curved (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015). By simulation, we also numerically determined the impact of finite dictionary size, by removing the rare CRCPD types from consideration: the type-token plot becomes more curved with a reduction of dictionary size. The previous attempts to derive the type-token relation, with finite or infinite dictionary size do not present the result in closed-form analytic expression (Eliazar, 2011; Lü et al., 2013; Tunnicliffe and Hunter, 2022). Our quadratic regression provides a simple formula for an empirical description of the systematic deviation from the Heaps’ law.
The definition of “type” or “dictionary word”, or in our case, Pfam domains, is supposed to be given by the database, and synonyms are not merged. For example, there are 13 protein domain names containing the prefix of zf-C2H2, including the top ranking zf-C2H2 (5493 appearances in the genome) and second ranking zf-C2H2_6 (1609 appearance in the genome). The reason of not combining potential synonyms is that it should be a feature of a “dictionary”, and actually, Heaps’ law partially reflects this lexical diversity. If one would combine CRCPD that are “similar” (up to certain degree), the type-token plot and Heaps’ law in human languages should also be modified accordingly.
As argued in (Lynch and Conery, 2003), a consequence of smaller population size is that selection pressure is much less effective in removing individuals without a perfect fit for the given environment. As nonadaptive forces dominated, genomes became less compact, with more redundancies, and inefficiencies were tolerated. On the other hand, this near-neutral random walk in Wright’s landscape allows a species to reach a new adaptive peak, potentially higher than the previous one (Wagner, 2019). Microorganisms have populations several orders of magnitude larger than animals. The higher level of genome redundancy and complexity in animals, versus the lower level of genome complexity in microorganisms, is then expected. For a genome without redundancy, the number of token is roughly equal to the number of types, and Heaps exponent is one. With redundancy, exponent is expected to be less than 1.
It is suggested in (Caetano-Anollés et al., 2017; Caetano-Anollés, 2021) that viral, archaeal and bacterial proteomics have a higher level of vocabulary growth than eukarya, as indicated by the higher value of $\alpha$ exponents (around 0.8). The $\alpha$ exponent for eukarya genomes in (Caetano-Anollés, 2021) is only around 0.1, in contrast with our exponent $\alpha\sim 0.7$ within one genome. It is tempting to use the slope in linear regression to rank different animal genomes in their level of duplication of protein domains. However, the limited x-range makes the estimation of slope less reliable. Also, when x-range is expanded, the linear (i.e., power-law when x and y are not in log-scale) trend becomes quadratic, and there are three parameters to characterize the type-token function, not just one. The similarity of all three fitting parameters for the animal genomes in Fig.7 make them more similar than different.
[…]
Despite the deviation from a perfect power-law for type-token plot, or in an alternative wording, that the power-law Heaps’ law is only technically correct in a limited range, there are other important lessons in this study from the linguistic perspective. Just like human languages with hundreds of new English words being created annually, albeit some of them are synonyms to the existing words, new protein domains have been generated during evolution. Even if the new protein domains may not be completely new, being at various degrees similar to the old ones, these may not be that different from the new synonymous words in human languages. Some of the seemingly non-novel new protein domains may continue to evolve and eventually lead to a new meaning and a new structure.
Acknowledgment
We would like to thank Patrick Villanueva for participating to the initial stage of the project, Tom MacCarthy and Roman Samulyak for useful suggestions.